OpenStudy (anonymous):

Is 3 5 9 15 ... an AP? If yes what's the common difference?

OpenStudy (radar):

2, 4, 6 2n

OpenStudy (anonymous):

@radar 2n is the common difference?

OpenStudy (anonymous):

So it is an ap?

OpenStudy (amistre64):

does 3-5 = 5-9 ?

OpenStudy (amistre64):

if the difference betwen each successive term is the same, then it has a "common" difference

OpenStudy (anonymous):

Ohkay. But we can see it is a sequence right ?

OpenStudy (amistre64):

it is a sequence, but it is not AP or GP; but we can devise a rule for it nonetheless

OpenStudy (amistre64):

i would have to dbl chk, but im thinking a sequance of numbers need not even have a rule to define it; but the definable ones are most useful to us

OpenStudy (anonymous):

You 're always a great help.

OpenStudy (amistre64):

its been known to happen :)

OpenStudy (anonymous):

One last question: sequence need not be an AP or a GP necessarily right ? As i can conclude

OpenStudy (anonymous):

Correct, this one appears to be n^2-n+3

OpenStudy (radar):

For the sequence at 15 which was n=4 or the 4th term the next "common difference" would of been 5(2) or 10 giving you 25 for the next term. However, as pointed out the term "common difference" does not apply, as the difference keeps growing as the sequence progresses, and estudier as provided an equation where if you know the sequence number you can predict the value, example for n=5, 25-5+5=25. That equation captures the progression of the sequence.